# Einstein’s basement - a model for dark matter and an expanding universe?

### v1

@@ Abstract
We present a model treating the energy-momentum relation of relativistic particles as the upper branch of a generalized energy-momentum relation emerging from a forbidden crossing between the constant energy of a massive particle and the photon line. The lower branch, a regime dubbed as Einstein’s basement, gives rise to particles with different kinematics that is analysed in the low-velocity limit. Allowing for gravitational interaction between those particles, and between particles of both branches, we discuss whether particles in Einstein’s basement might be suitable dark matter candidates. Moreover, the special dynamics of mixed systems leads to an expansion mechanism for regular matter in space. Tests of the presented model by astronomical observations are suggested.
Introduction
The large scale dynamics of the Universe by regular matter is understood only if we allow for hypothetical dark matter and dark energy [1]. Dark matter [2] is a well established concept, allowing to describe a huge number of observations in cosmology [3, 4], besides other explanations such as “Modified Newtonian Dynamics (MOND)” [5, 6]. Whereas regular matter is successfully described by the standard model of particles [7], the origin of dark matter is not yet clear. A number of theoretical hypotheses has been offered for the origin of dark matter like primordial black holes [8], weakly interacting massive particles (WIMPs) [9], axion-like particles [10], massive compact halo objects (MACHOs) [11], dark photons [12], erebons [13], Kaluza-Klein particles [14] or others. Another attempt to explain dark matter and dark energy was based on the assumption of negative masses [15]. More than fifty direct or indirect experimental approaches for detection of the WIMPs and axions are under way [16] but up to now none of the proposed candidate particles have been found to explain dark matter to the extent to which it is present in the universe [17, 18].
Any attempt to solve the presently unexplained mysteries is stepping back from the known physics in one point or another. A systematic proof of the assumptions of standard model physics therefore is essential for constraining New Physics, and the degree of how strong known principles of physics can be violated. Here, we explore a possible extension where we will relax our restrictions on what is considered to be a particle.
We rely on a more general view of a “particle” as any entity with a well defined energy momentum dispersion relation (EMR). Among those are also so-called “quasi particles” or “collective excitations” where a particular interaction dresses the bare particles with modified properties. This fruitful concept is used, e.g., in Fermi liquids [19], phonons or polaritons [20, 21, 22] in solid state physics, but may also apply to the Standard Model where particles, normally considered as fundamental, are treated as excitations of their underlying quantum fields.
The paper is organized as follows. First, we find besides the well known relativistic EMR a second one whose energy is below the former, marking a region that is dubbed henceforth as “Einstein’s basement”. Then, as a logical first step, we apply classical mechanics to the new EMR and derive the corresponding equations of motion analogue to the mechanics of special relativity and Newtonian physics. We then explore whether such quasi particles might represent candidates for dark matter and give examples of possibly observable phenomena.
In special relativity a free particle has the well-known EMR
$E_{SR}(p)=\sqrt{(mc^{2})^{2}+{\mathbf{p}}^{2}c^{2}}$ (1)
@@ Introduction
with $m$ the rest mass of the particle, $c$ the speed of light in vacuum and $p=|{\mathbf{p}}|$ the modulus of the momentum of the particle [23, 24]. This EMR is (the only one) intimately associated with the geometry of Minkowski spacetime governed by Lorentz symmetry. Its properties have been tested to its best over the last century. This EMR is shown in Fig. 1 as the upper two curves for two different masses $m_{1}$ and $m_{2}$. For low velocities each curve starts with the respective rest mass energy $m_{i}c^{2}$ before it turns into a linear behavior for large kinetic energies.
I The model
@@
Our model starts with the notion that the EMR of special relativity represents the upper branch of an avoided crossing between the EMR of a constant matter bare state $E_{b1}=mc^{2}$ and the EMR of a photon bare state $E_{b2}=pc$. In a simple form an avoided crossing can be described by a matrix $\cal{A}$ that couples these two bare states with the coupling $F(p)$ as
@@
Figure 1: The relativistic energy momentum relations of two free particles (upper two curves) with rest masses $m_{1},m_{2}$ as function of the particles’ momenta. They approach the respective rest energies $E=m_{i}c^{2}$ (horizontal lines) for low momenta and the energy-momentum relation of a photon (dash-dotted line) for high momenta. Furthermore, for each $m_{i}$ a second branch is shown, which, in contrast to the preceding case, asymptotically approaches the photonic line at low momenta and the respective mass lines at high momenta. The gray area represents what we refer to as Einstein’s basement in this paper.
The new EMR of Eq. (4) is displayed also in Fig. 1 for the same two masses $m_{1},m_{2}$. For small momenta, they approach the photon line whereas they cling to their rest masses for large momenta.
[…]
@@
Thus, we find that the velocity of a particle behaves differently in both cases. In special relativity $|\mathbf{v}|=v$ depends linearly on the momentum for $v\ll c$ and approaches the velocity of light for asymptotically high momenta. In Einstein’s basement $v$ decreases from $c$ for low momenta and goes to zero for $p\to\infty$. Both behaviors are illustrated in Fig. 2.
@@
We find that both momenta have a similar form, while in Eq. (9) the velocity $v$ is replaced by $c-v$ in comparison to the relativistic case. Thus, we observe that both branches respect $c$ as an ultimate limit for the motion of particles, while the roles of $v=0$ and $v=c$ are interchanged. In the relativistic case, the special relation of the momentum (8) and the energy (1), gives rise to the class of Lorentz transformations, keeping $\sqrt{E_{SR}^{2}-c^{2}\mathbf{p}_{SR}^{2}}=mc^{2}$ invariant. However, this does not hold for the lower branch, where no such transformation can be found. In particular, $\sqrt{E_{EB}^{2}-c^{2}\mathbf{p}_{EB}^{2}}$ is not Lorentz invariant by itself, i.e., has a different value in every Lorentz frame, giving up this principle of relativity in the lower branch solely. We will see that this feature will also translate into the equations of motion, when we investigate the dynamics of the newly introduced particles.
To study the kinematics of particles in Einstein’s basement, we want to construct the counterpart of the well known relativistic Lagrangian
@@
for a free particle in Einstein’s basement, where the integration constant was chosen as $-{m}c^{2}$ to obtain the Lagrangian for a particle at rest. In analogy to the second line of Eq. (I), we obtain a kinetic energy term in the low-velocity limit. Remarkably, also in this limit, the speed of light $c$ remains as a parameter in the system, affecting the motion of such particles. In the next section, we will analyze this expression and discuss the resulting dynamics of particles, which qualitatively differs from the Newtonian case in many aspects.
II New dynamics in Einstein’s basement
In the following, we have a closer look at the dynamics in Einstein’s basement by using the free particle Lagrangian Eq. (I) in the low-velocity limit and complement it by an additional potential depending on the particle’s position in space
@@ II New dynamics in Einstein’s basement
where $\mathbf{a}=\ddot{\mathbf{q}}$ is the particle’s acceleration and $\mathbf{F}$ the force resulting from the potential. We find that the time derivative of the momentum $\frac{d}{dt}\mathbf{p}_{EB}\not=m\,\mathbf{a}$, does not coincide with the acceleration, as it is the case in Newtonian physics. In contrast, $m\,\mathbf{a}$ is complemented by an additional term, which depends on the particle’s velocity $\mathbf{v}$. By rearranging Eq. (15) in the form
@@
II.1 Dynamics of two classical particles
@@
II.2 Dynamics of two slow basement particles
@@
of two gravitationally interacting particles using Eq. (I) in the low-velocity limit. The equation of relative motion for the two particles reads
@@
II.3 Dynamics of a classical and a (slow) basement particle
@@
Figure 4: Simulated trajectories of a regular particle with mass $m_{2}$ (black) and a slow basement particle with mass $m_{1}$ (blue), interacting gravitationally. The basement particle is almost unaffected by its companion, whereas the regular particle follows the former one in a spiraling motion, see (a) and (b). The relative motion of $m_{2}$ with respect to $m_{1}$, given in (c), shows two periodicities: an elliptic orbit comparable to Kepler motion, and an oscillation of the apsidal position.
III Are particles in Einstein’s basement dark matter candidates?
We are now in the position to investigate whether our model can reproduce the central features of dark matter.
The mass of dark matter.
In our approach we treat the EMR of regular matter as one of the two branches of a generalized energy-momentum relation that emerges from a forbidden crossing of the constant energy of a massive particle and the photon line. This leads to a new zoo of particles, having the same masses as their relativistic counterparts. The emerging new sector could complement the known constituents of the current standard model.
The existence of dark matter was first deduced by Fritz Zwicky [30] from observed discrepancies in the dynamics of galaxy clusters from the laws of celestial mechanics, such as the virial theorem. In contrast to the usual approach to attribute these observations solely to an invisible distribution of matter with regular kinematics, in our model the emerging particles have a different kinematics as in the regular relativistic or Newtonian case. This mechanism can be an alternative to the usual dark matter model.
Photons, and why the basement may be dark.
There is a striking difference between the regular EMR $E_{SR}(p)$ and $E_{EB}(p)$ in Einstein’s basement with respect to the possibility of mass-less particles by setting $mc^{2}=0$. In the first case we recover $E_{SR}(p)=pc$, i.e., the linear EMR of photons. In contrast, for a mass-less particle in Einstein’s basement we find $E_{EB}(p)=0$, such that it never has an energy other than zero. Consequently, there are no photon-like particles in Einstein’s basement. This, a priori, would not prevent that a particle in Einstein’s basement could interact with a regular photon. The interaction between a photon and a particle in Einstein’s basement would require an additional coupling between both branches. Let us look on the probability of such a hypothetical interaction from a different viewpoint. Diabatic electromagnetic transitions are well known to occur in the case of an avoided crossing. Such transitions were modeled long ago by Landau [31], Zener [32], Stückelberg [33] and Majorana [34] and have been tested experimentally in quantum optics [35]. Assuming that diabatic transitions under the influence of one or more of the well known interactions are allowed, they would lead to an increased population in the $E_{EB}$ state which has a lower energy than $E_{SR}$. Consequently, direct transitions between the two regimes, and, hence, the direct interaction between both states must be extremely low. As a result, we conclude that there is not much room for the interaction of photons with particles in Einstein’s basement.
Low coupling and the predominance of dark matter over regular matter
While we can assume that a spontaneous decay of classical particles into the basement is highly suppressed at the current age of the universe, the interaction between both branches may have changed during its different phases. However, the basement as the energetically more favorable state can be an explanation of the current dominance of dark matter in comparison to common matter in the present universe.
Clusters of dark matter.
For regular matter, the formation of large bodies by accumulation of microscopic particles via gravitational attraction is well understood. A possible aggregation of dark matter consisting of particles from Einstein’s basement would be more intricate due to the repulsive and attractive forces in the game.
The velocities of cold dark matter in galaxies and spiral nebula have been determined to be in the range of $200-550$ km/s [36, 37]. As we discussed in Sec. II.3, regular matter is strongly attracted by such slow basement particles, see Eq. (27), while the potentially repulsive forces between the basement particles are suppressed by $(v/c)^{3/2}$, see Eq. (24). Moreover, given the attractive forces between regular matter components, such a mixture could be stable even though the fast constituents of dark matter would boil off.
The preceding arguments lead us to assume that the quasi particles derived from our model of Einstein’s basement can be regarded as possible dark matter candidates.
Having discussed the relation of basement particles and the observed properties of dark matter, we want to emphasize yet another feature of a mixed system of basement particles and regular matter. Under the assumption of classical gravitational interaction, particles in Einstein’s basement - apart from special cases - are repelled by each other, while regular particles follow them regardless of their own mass. This mechanism describes an expansion of a mixed system in space. Whether this can be a source of the observed accelerated expansion of the universe, as it is suggested by the study of type Ia supernovae candles [38, 39], has to be investigated. The modeling of statistically significant ensembles is indispensable to answer this question. It is also necessary to check if our model can explain other observations already described reasonably well by the models of $\Lambda$CDM and MOND at large and small scales, respectively.
IV Summary, Outlook and Conclusions
[…]
Our findings might possibly be tested quantitatively at different scales, for instance by astrophysical observations of stars, nebulae and galaxies and by gravitational lensing effects of regular and dark matter components. A promising candidate for such an investigation could be the so-called bullet cluster [40]. This cluster comprises of two components each with a spearhead dark matter contribution followed by a distribution of hot regular gaseous material. In such a geometry our model asks - in contrast to regular kinematics - for additional acceleration of the two dark matter components away from each other, which might be observable in successive measurements.
It must be checked, whether our model delivers an competitive explanation for the existing measurement data and provides an alternative to the current maps of the dark matter distribution in the galaxies [41, 42]. Revisiting the huge number of existing data and the new data expected from the EUCLID space mission [43] could give a hint whether the presented ideas should be pursued further.

### v2

# Einstein’s basement - A kinematic sector complementing special relativity

@@ Abstract
We revisit the concept of particles as it is used in special relativity. The presented model treats the energy-momentum relation of relativistic particles as the upper branch of a generalized energy-momentum relation of quasi particles. These particles emerge from a forbidden crossing between the constant energy of a massive particle in rest and the linear energy momentum relation of a massless particle. The lower branch, a regime dubbed as Einstein’s basement, gives rise to particles with different kinematics that is analyzed in the low-velocity limit. Allowing for gravitational interaction between those particles we find both attraction and repulsion, depending on their velocity with respect to an absolute space. This absolute frame is only relevant for Einstein’s basement and does not affect the relativistic dynamics of regular matter as long as no coupling between the branches is considered, while weak coupling induces local Lorentz violations. We finally discuss whether our approach can be used to model phenomena of dark matter and dark energy.
I Introduction
In classical mechanics, the concept of a particle has proven extremely useful to abstract the motion of isolated bodies. This concept of particles has the advantage that almost everybody has a clear intuition about its meaning. However, the interpretation of what a particle is becomes more complicated when it goes to the physics of the smallest scales described within the framework of quantum mechanics and ultimately by quantum field theory [1]. There, “particles” are the excitations of quantum fields.
This picture holds not only for fundamental fields, and the corresponding fundamental particles, such as electrons and photons, but also can be used to describe emergent phenomena, e.g. in solid state physics, where the excitations of non-fundamental fields are treated as quasi particles. Fundamental particles differ from quasi particles like phonons or polaritons [2, 3, 4] in the fact that no underlying system they emerge from, such as an atomic lattice or a polarizable optical medium, is known. However, history tells us that the term “fundamental” always shifts to a more basic layer when our understanding of the world and our insight into the underlying mechanisms increases. While once the atom was considered as indivisible, in the last century one discovered protons and neutrons and even found that they are compounds of quarks and gluons as the particles of quantum chromodynamics [5], considered as fundamental to date.
Quasi particles, as well as fundamental particles have a well defined energy momentum relationship (EMR), emerging from underlying symmetry principles. A prominent example is the standard model of particles [6] where the irreducible representations of the Lorentz symmetry group give rise to the zoo of particles, respecting the EMR of special relativity [7]. While being the most efficient description of nature at small scales we have today, the standard model still is capable to explain a number of observations [8]. Second, the dynamics of regular matter in galaxies hints on the presence of additional, dark matter [9, 10, 11] built up by yet unknown particles.
One way to tackle these problems might be to relax the restrictions on what is considered to be a “fundamental” particle and, instead, to assume such particles to be quasi particles of a yet unknown underlying system.
Following this line, our paper is organized as follows. In Sec. II, we apply the no-level crossing theorem to the two conceptually different types of particles, namely those with constant rest energy and those with purely kinetic energy. Below the well known EMR of special relativity we obtain a second one, marking a region that is dubbed henceforth as “Einstein’s basement”. We then discuss the formal equivalence of the derived energy momentum relationships with the ones [2, 3, 4] encountered in the polariton system (Sec. III). In Sec. IV we apply classical mechanics to the new EMR and derive and discuss the corresponding equations of motion analogously to the mechanics of special relativity and Newtonian physics. We then study two-particle systems of gravitationally interacting quasi particles from both regions and discuss consequences of our findings and possible extensions of our model in Sec. V.
Focus of this work is to consider the dynamics of classical point particles with a modified dispersion relationship in comparison to Newtonian mechanics. These studies lay the foundation to investigate further concepts like spin or gauge invariance that could be considered when fields and their dynamics are constructed “on top” of the given energy-momentum relationship, and are quantized afterwards.
II The model
Our model starts with the notion of two conceptually different types of bare particles, namely those with a constant rest energy $E=mc^{2}$ and those with only kinetic energy $pc$ (e.g. light). Here, $E$ denotes the energy, $m$ the rest mass, and the $c$ the speed of light in vacuum, respectively. $p=|{\mathbf{p}}|$ is the modulus of the momentum of the particle [12, 13]. We see from Fig. 1 that these two EMRs (i.e. particles) have the same energy for $p=mc$, which means that their EMRs have an allowed crossing. However, for any coupling (interaction) between these particles the crossing will be avoided. It is a universal principle of physics - sometimes called no-level crossing theorem - that a pair of energy levels connected by a perturbation do not cross as the strength of the perturbation is varied from zero [14, 15]. Examples of avoided crossings are found everywhere in macroscopic physics and quantum mechanics, e.g. in mechanics [16, 17, 18], optics [3], acoustics [19], in quantum mechanics of molecules [20] or quantum chemistry [21], to name only a few. In the avoided crossing we have now two branches of the EMRs.
In our model we identify the upper branch with the well-known EMR of special relativity
$E_{SR}(p)=\sqrt{(mc^{2})^{2}+{\mathbf{p}}^{2}c^{2}}.$ (1)
@@ II The model
This EMR is (the only one) intimately associated with the geometry of Minkowski spacetime governed by Lorentz symmetry. Its properties have been tested to its best over the last century. This EMR is shown in Fig. 1 as the upper curve for the mass $m$. For low velocities it starts with the rest energy $mc^{2}$ before it approaches a linear behavior for large kinetic energies.
@@
In a simple form, an avoided crossing can be described by a matrix $\cal{A}$ that couples the two bare states with the coupling $F(p)$ as
@@
Figure 1: The energy momentum relation (EMR) of a particle with constant rest energy (full line) and the EMR of a particle with pure kinetic energy $pc$ (dashed line) cross at the energy $E=mc^{2}$ (red dot). In our model the coupling of these states leads to an avoided crossing with the EMR of a free particle according to special relativity (dotted line) and the EMR of a particle in Einsteins basement (blue area; dashed-dotted line).
The new EMR of Eq. (4) is displayed also in Fig. 1 for the same mass $m$. For small momenta, the EMR approaches the $pc$-line whereas it clings to the rest mass for large momenta. With the coupling $F(p,m)$, the EMRs $E_{SR}$ and $E_{EB}$ represent quasi particles that can be regarded as the excitations of an (up to now) unknown field.
In what follows, we will investigate how the kinematics of the particles can change when the physics of the lower branch becomes accessible.
@@
Thus, we find that the velocity of a particle behaves differently in both cases. In special relativity $|\mathbf{v}|=v$ depends linearly on the momentum for $v\ll c$ and approaches the velocity of light for asymptotically high momenta. For particles in Einstein’s basement $v$ decreases from $c$ for low momenta and goes to zero for $p\to\infty$. Both behaviors are illustrated in Fig. 2.
@@
In Einstein’s basement, the well known Lorentz factor of special relativity
$\gamma_{SR}=\frac{1}{\sqrt{1-(\frac{v}{c})^{2}}}$ (10)
is replaced by a corresponding factor
$\gamma_{EB}=\frac{1}{\sqrt{1-(\frac{c-v}{c})^{2}}}.$ (11)
We find that both momenta have a similar form, while in Eq. (9) the velocity $v$ is replaced by $c-v$ in comparison to the relativistic case [25].
Hence, both, (10) and (11) are linked by a mirror symmetry [25]
$v\leftrightarrow c-v$ (12)
which is visualized in Fig. 3.
Figure 3: In velocity space the constant speed of light is visualized by a sphere where a particular velocity $\mathbf{v}$ and the corresponding $\mathbf{c}-\mathbf{v}$ are depicted. We see that for each velocity $\mathbf{v}$ one can find a parallel vector ${\mathbf{c}}$ to construct $\mathbf{c}-\mathbf{v}$ with a magnitude $c-v\in[0,c]$, used in Eqs. (9) and (11).
Thus, we observe that both branches respect $c$ as an ultimate limit for the motion of particles, while the roles of $v=0$ and $v=c$ are interchanged. In the relativistic case, the special relation of the momentum (8) and the energy (1), gives rise to the class of Lorentz transformations, keeping $\sqrt{E_{SR}^{2}-c^{2}\mathbf{p}_{SR}^{2}}=mc^{2}$ invariant. However, this does not hold for the lower branch, where no such transformation can be found. In particular, $\sqrt{E_{EB}^{2}-c^{2}\mathbf{p}_{EB}^{2}}$ is not Lorentz invariant by itself, i.e., has a different value in every Lorentz frame, giving up this principle of relativity in the lower branch. We will see that this feature will also translate into the equations of motion, when we investigate the dynamics of the newly introduced quasi particles.
Group theory [26] and differential geometry [27] allow us to get deeper insights in the framework behind special relativity and Einstein’s basement. With the mirror transformation Eq. (12) (i.e. an opposite isometry) between the dynamics of both sectors, an overarching symmetry group must complement the Poincaré group by shifts and reflections in velocity space. The full elaboration of these thoughts in group theory will not be treated here, since here we are primarily interested in the novel kinematics in Einstein’s basement to be discussed in Sec.  IV. But before, we will discuss a quantum optical system in solid state physics that already displays the same essential properties as our model does.
III Polaritons - a quantum optical system equivalent to our model
An equivalent system to Einstein’s basement can be found e.g. in the (phonon)-polaritons in an ionic crystal [2, 3, 4] or in the coupled magnon-phonon modes in a ferromagnet in the Faraday configuration [28]. The phonon-polariton system shows the typical dispersion relations of Fig. 4.
Figure 4: The energy momentum relationships of phonon-polaritons after [3] (see text).
In the ionic crystal usually two mechanical eigenfrequencies can be excited that correspond to transverse and longitudinal oscillations (phonon modes) [29] of the atomic layers [2]. An electromagnetic wave can interact with the ions via its electric field. The EMRs of this system (Fig. 4) resemble strikingly the EMRs of Fig. 1, which is not surprising since one starts with the same bare states of light and modes with constant energies that we have used to construct our model. In this system we find a gap resulting from the different frequencies of the longitudinal and transversal modes. This gap is not present in model (Fig. 1) as it is also the case in some polariton systems, such as in germanium or silicon crystals [2]. Observers in this system who are aware of the upper branch only, would conclude (like we do in special relativity) that Lorentz transformations are valid everywhere and that the physical laws have to be the same in each inertial system. If they would find the lower branch, however, these observers would conclude that both sorts of EMRs i.e. both sorts of quasi particles can persist in the same physical system as we investigate in this paper.
Such a “pseudo-relativistic behavior” of the upper EMR of a polariton system has also been observed in other crystalline systems [30, 31]. The formally identical conditions between our model and the polariton system in a crystal could be used as a toy model for experiments to visualize special properties of particles in Einstein’s basement and their interaction [32].
IV New dynamics for particles from Einstein’s basement
To study the kinematics of particles of Einstein’s basement, we want to construct the counterpart of the relativistic Lagrangian
@@ IV New dynamics for particles from Einstein’s basement
for a free particle of Einstein’s basement, where the integration constant was chosen as $-{m}c^{2}$ to obtain the Lagrangian for a particle at rest. In analogy to the second line of Eq. (IV), we obtain a kinetic energy term in the low-velocity limit of Eq. (IV). Remarkably, also in this limit, the speed of light $c$ remains as a parameter in the system, affecting the motion of such particles. In the next section, we will analyze this expression and discuss the resulting dynamics of particles, which qualitatively differs from the Newtonian case in many aspects.
In the following, we have a closer look at the dynamics in Einstein’s basement by using the free particle Lagrangian Eq. (IV) in the low-velocity limit and complement it by an additional potential depending on the particle’s position in space
@@
where $\mathbf{a}=\ddot{\mathbf{q}}$ is the particle’s acceleration and $\mathbf{F}$ the force resulting from the potential. We find that the time derivative of the momentum $\frac{d}{dt}\mathbf{p}_{EB}\not=m\,\mathbf{a}$, does not coincide with the acceleration, as it is the case in Newtonian physics. In contrast, $m\,\mathbf{a}$ is complemented by an additional term, which depends on the particle’s velocity $\mathbf{v}$. By rearranging Eq. (18) into the form
@@
IV.1 Dynamics of two classical particles
@@
IV.2 Dynamics of two slow basement particles
@@
of two gravitationally interacting particles using Eq. (IV) in the low-velocity limit. The equation of relative motion for the two particles reads
@@
IV.3 Dynamics of a classical and a (slow) basement particle
@@
Figure 6: Simulated trajectories of a regular particle with mass $m_{1}$ (black) and a slow basement particle with mass $m_{2}$ (blue), interacting gravitationally. The basement particle is almost unaffected by its companion, whereas the regular particle follows the former one in a spiraling motion, see (a) and (b). The relative orbit of $m_{1}$ with respect to $m_{2}$, given in (c), shows two periodicities: an elliptic orbit comparable to Kepler motion, and an oscillation of the apsidal position [35].
V Summary, Conclusions and Outlook